At what speed is a car approaching a
stationary observer if he hears the music of
the car radio with a frequency 10% higher
than the actual frequency ? Speed of
sound =340 ms^-1.
1
Expert's answer
2019-05-21T08:54:25-0400

Doppler effect is given by formula

"\frac{\Delta f}{f}=\frac{v}{c}"

So, the speed of a car

"v=c\frac{\Delta f}{f}"

"=340\:\rm{m/s}\times 0.1=34\:\rm{m/s}"